A spinner game has a wheel with the numbers zero through 28 marked in equally spaced slots. Consider a $1 bet on a number from 1 to 28. If your number hits, you have $28 returned to you; otherwise, you lose your $1 bet. Find the expected return for this bet. (Round your answer to two decimal places.)
Answer #1

P(winning $28) = 1/28

P(loosing $1) = 27/28

Expected return for this bet = 28 * 1/28 - 1 * 27/28 = $0.04 (ans)